# Large Scale Diffusion Distillation via Score-Regularized Continuous-Time Consistency

### v1

@@ 4.1 Quality Repair with Score Regularization
As shown in Figure 4, we mitigate quality limitations of sCM by introducing score-based regularization on long-skip consistency, which complements sCM with reverse divergence. While SiD (Zhou et al., 2024) achieves impressive results on academic benchmarks, we observe no clear advantage in T2I and T2V tasks and instead adopt the more memory-efficient DMD (Yin et al., 2024b):
@@
where $\bm{f}_{\text{fake}}$ is the denoiser of the fake score network, $p_{D}$ is a time distribution and sg is the stop-gradient operator. The final rCM objective is $\mathcal{L}_{\text{rCM}}(\theta)=\mathcal{L}_{\text{sCM}}(\theta)+\lambda\mathcal{L}_{\text{DMD}}(\theta)$, where $\lambda$ is a balancing weight. Empirically, we find $\lambda=0.01$ generalizes across all models and tasks.
@@ 4.2 Stable Time Derivative Calculation
High-Precision Time. Finite-difference approximation, however, is sensitive to $\Delta t$ and becomes unstable for 10B+ models and video tasks. In such cases, we enforce FP32 precision for all time embedding layers using the torch.amp.autocast context, as in Wan. Although this introduces an initial mismatch with pretrained Cosmos networks, it ensures stable rCM training.
@@ 5.1 Experimental Setups
Models, Tasks and Datasets. To demonstrate scalability and performance of rCM, we distill Cosmos-Predict2 (NVIDIA, 2025) T2I models (0.6B, 2B, 14B) and Wan2.1 (Wan et al., 2025) T2V models (1.3B, 14B). We leverage curated data from NVIDIA (2025), supplemented with synthetic data generated by Wan2.1 T2V 14B for Wan distillation. In principle, the training could also rely solely on teacher-generated synthetic data, as in Yin et al. (2024b; 2025); Huang et al. (2025).
@@ 5.2 Results
Cosmos-Predict2 TI2V (NVIDIA, 2025) 2B $1280\times 704\times 93$ $35\times 2$ 0.32 83.03 88.6
@@ C.2 Network Restructuring
To make JVP computation compatible with Fully Sharded Data Parallel (FSDP), we restructure the forward functions of network layers. Specifically, we define a base class JVP (Listing 7) that extends torch.nn.Module and supports both standard forward execution and JVP-mode execution. When withT=True, the forward pass receives and returns both the primals and their tangents, with each primal and the correpsonding tangent wrapped in the TensorWithT tuple type.
@@ Appendix D Experiment Details
$p_{G}$
$z\sim\mathcal{N}(-0.8,1.6)$
$t=\arctan(z)$
$z\sim\mathcal{N}(-0.8,1.6)$
$t=\arctan(z)$
$z\sim\mathcal{N}(-0.8,1.6)$
$t=\arctan(z)$
$z\sim\mathcal{N}(-0.8,1.6)$
$t=\arctan(\sqrt{T}z)$
$z\sim\mathcal{N}(-0.8,1.6)$
$t=\arctan(\sqrt{T}z)$
$p_{D}$
$z\sim\mathcal{N}(0.0,1.6)$
$t=\arctan(z)$
$z\sim\mathcal{N}(0.0,1.6)$
$t=\arctan(z)$
$z\sim\mathcal{N}(0.0,1.6)$
$t=\arctan(z)$
$u\sim\mathcal{U}(0,1)$
$t_{\text{RF}}=\frac{5u}{1+4u}$
$t=\arctan\left(\frac{t_{\text{RF}}}{1-t_{\text{RF}}}\right)$
$u\sim\mathcal{U}(0,1)$
$t_{\text{RF}}=\frac{5u}{1+4u}$
$t=\arctan\left(\frac{t_{\text{RF}}}{1-t_{\text{RF}}}\right)$
@@ Appendix F Prompts
Animated scene features a close-up of a short fluffy monster kneeling beside a melting red candle. The art style is 3D and realistic, with a focus on lighting and texture. The mood of the painting is one of wonder and curiosity, as the monster gazes at the flame with wide eyes and open mouth. Its pose and expression convey a sense of innocence and playfulness, as if it is exploring the world around it for the first time. The use of warm colors and dramatic lighting further enhances the cozy atmosphere of the image. Figure 1

### v2

@@ 4.1 Quality Repair with Score Regularization
As shown in Figure 4, we mitigate quality limitations of sCM by introducing score-based regularization on long-skip consistency, which complements sCM with reverse divergence. While SiD (Zhou et al., 2024) achieves impressive results on academic benchmarks, we observe no clear advantage in T2I and T2V tasks (Figure 1) and instead adopt the more memory-efficient DMD (Yin et al., 2024b):
@@
where $\bm{f}_{\text{fake}}$ is the denoiser of the fake score network, $p_{D}$ is a time distribution and sg is the stop-gradient operator. The final rCM objective is $\mathcal{L}_{\text{rCM}}(\theta)=\mathcal{L}_{\text{sCM}}(\theta)+\lambda\mathcal{L}_{\text{DMD}}(\theta)$, where $\lambda$ is a balancing weight. Empirically, we find $\lambda=0.01$ generalizes across our used models and tasks.
@@ 4.2 Stable Time Derivative Calculation
High-Precision Time. Finite-difference approximation, however, is sensitive to $\Delta t$ and becomes unstable for 10B+ models and video tasks. In these regimes, we revert to the native continuous-time derivative computation via full JVP, but enforce FP32 precision for all time embedding layers using the torch.amp.autocast context (as done in Wan). Although this introduces an initial mismatch with pretrained Cosmos networks, it ensures stable rCM training.
@@ 5.1 Experimental Setups
Models, Tasks and Datasets. To demonstrate scalability and performance of rCM, we distill Cosmos-Predict2 (Ali et al., 2025) T2I models (0.6B, 2B, 14B) and Wan2.1 (Wan et al., 2025) T2V models (1.3B, 14B). We leverage curated data from Ali et al. (2025), supplemented with synthetic data generated by Wan2.1 T2V 14B for Wan distillation. In principle, the training could also rely solely on teacher-generated synthetic data, as in Yin et al. (2024b; 2025); Huang et al. (2025).
@@ 5.2 Results
Cosmos-Predict2 TI2V (Ali et al., 2025) 2B $1280\times 704\times 93$ $35\times 2$ 0.32 83.03 88.6
@@
Figure 7: Video samples from 4-step Wan2.1 1.3B rCM models under different $\lambda$. For each prompt, we use 5 different random seeds to demonstrate diversity. VBench scores are in the parentheses.
Ablation Study on $\lambda$. In principle, the balancing weighting $\lambda$ between sCM and DMD losses should control the trade-off between diversity (mode-covering) and quality (mode-seeking). To validate this, we perform a grid search over $\lambda\in\{1,0.1,0.01,0.001\}$ on Wan2.1-1.3B, training each model with a batch size of 64 for 10k iterations. As shown in Figure 7, larger $\lambda$ (i.e., stronger DMD weighting) results in better quality but less diversity, while smaller values exhibit the opposite trend. At a granularity of one order of magnitude, we find that $\lambda=0.01$, as the smallest scale to preserve good quality, offers a “sweet spot” balancing both quality and diversity.
@@ 6 Conclusion
The Use of Large Language Models (LLMs)
We used large language models (LLMs) solely as a writing assistant for language polishing and improving clarity of presentation. The LLMs were not involved in research ideation, methodological design, experimental execution, or result analysis. All scientific contributions and substantive writing were carried out by the authors.
@@ C.2 Network Restructuring
To make JVP computation compatible with Fully Sharded Data Parallel (FSDP), we restructure the forward functions of network layers. Specifically, we define a base class JVP (Listing 8) that extends torch.nn.Module and supports both standard forward execution and JVP-mode execution. When withT=True, the forward pass receives and returns both the primals and their tangents, with each primal and the correpsonding tangent wrapped in the TensorWithT tuple type.
@@ Appendix D Experiment Details
$p_{G}$ $\log z\sim\mathcal{N}(-0.8,1.6^{2})$ $t=\arctan(z)$ $\log z\sim\mathcal{N}(-0.8,1.6^{2})$ $t=\arctan(z)$ $\log z\sim\mathcal{N}(-0.8,1.6^{2})$ $t=\arctan(z)$ $\log z\sim\mathcal{N}(-0.8,1.6^{2})$ $t=\arctan(\sqrt{T}z)$ $\log z\sim\mathcal{N}(-0.8,1.6^{2})$ $t=\arctan(\sqrt{T}z)$
$p_{D}$ $\log z\sim\mathcal{N}(0.0,1.6^{2})$ $t=\arctan(z)$ $\log z\sim\mathcal{N}(0.0,1.6^{2})$ $t=\arctan(z)$ $\log z\sim\mathcal{N}(0.0,1.6^{2})$ $t=\arctan(z)$ $u\sim\mathcal{U}(0,1)$ $t_{\text{RF}}=\frac{5u}{1+4u}$ $t=\arctan\left(\frac{t_{\text{RF}}}{1-t_{\text{RF}}}\right)$ $u\sim\mathcal{U}(0,1)$ $t_{\text{RF}}=\frac{5u}{1+4u}$ $t=\arctan\left(\frac{t_{\text{RF}}}{1-t_{\text{RF}}}\right)$
@@ Appendix F Prompts
Animated scene features a close-up of a short fluffy monster kneeling beside a melting red candle. The art style is 3D and realistic, with a focus on lighting and texture. The mood of the painting is one of wonder and curiosity, as the monster gazes at the flame with wide eyes and open mouth. Its pose and expression convey a sense of innocence and playfulness, as if it is exploring the world around it for the first time. The use of warm colors and dramatic lighting further enhances the cozy atmosphere of the image. Figure 1,7
The camera follows behind a white vintage SUV with a black roof rack as it speeds up a steep dirt road surrounded by pine trees on a steep mountain slope, dust kicks up from it’s tires, the sunlight shines on the SUV as it speeds along the dirt road, casting a warm glow over the scene. The dirt road curves gently into the distance, with no other cars or vehicles in sight. The trees on either side of the road are redwoods, with patches of greenery scattered throughout. The car is seen from the rear following the curve with ease, making it seem as if it is on a rugged drive through the rugged terrain. The dirt road itself is surrounded by steep hills and mountains, with a clear blue sky above with wispy clouds. Figure 7
A close up view of a glass sphere that has a zen garden within it. There is a small dwarf in the sphere who is raking the zen garden and creating patterns in the sand. Figure 7
A playful raccoon is seen playing an electronic guitar, strumming the strings with its front paws. The raccoon has distinctive black facial markings and a bushy tail. It sits comfortably on a small stool, its body slightly tilted as it focuses intently on the instrument. The setting is a cozy, dimly lit room with vintage posters on the walls, adding a retro vibe. The raccoon’s expressive eyes convey a sense of joy and concentration. Medium close-up shot, focusing on the raccoon’s face and hands interacting with the guitar. Figure 7